OpenStudy (anonymous):

According to my book, tan(-6pi) = tan 0 = 0. My calculator disagrees and gives tan(-6pi) = -0,34 Help?

OpenStudy (wikiemol):

Is your calculator in radian mode?

jimthompson5910 (jim_thompson5910):

make sure you're in radian mode

OpenStudy (anonymous):

You're both right, I wasn't in radian mode. But that just makes things weirder, not tan (-6*pi) = 2e^-13

OpenStudy (anonymous):

* "...weirder, now tan (-6*pi) = 2e^-13"

jimthompson5910 (jim_thompson5910):

well if you were you degree mode, then tan(-2pi) = -tan(2pi) tan(-2pi) = -tan(2*3.14) tan(-2pi) = -tan(6.28) tan(-2pi) = -0.11004772101689 So that's what you should have gotten IF you were you degree mode

jimthompson5910 (jim_thompson5910):

oh sry I did 2pi on accident, let me fix

jimthompson5910 (jim_thompson5910):

if you were you degree mode, then tan(-6pi) = -tan(2pi) tan(-6pi) = -tan(6*3.14) tan(-6pi) = -tan(18.84) tan(-6pi) = -0.3412069928076 So that's what you should have gotten IF you were you degree mode

OpenStudy (phi):

**now tan (-6*pi) = 2e^-13 2 times 10^-13 is a tiny number. It is not zero because your calculator does not have enough digits to do this problem accurately. But for practical purposes 2e-13 is close enough to zero.

OpenStudy (phi):

although my calculator comes up with 0.

OpenStudy (anonymous):

allright, I see, thanks!